A ball bearing similar to problem 1.1 has the foll
Answer #1

Calculate weight % of iron as follows: iron 00 100-sum of weight % (S + Si + P + Mn + C + Mo + Cr) = 100-(0.043 0.220 + 0.031

2. For Silicon, -weight % weight of ball 100 Weight- 0.220 x 3.756 g 100 = 8.2×10-3 g weight of element molar mass of element

4. For Manganese, Weight = weight % weight of ball 0.546 x 3.756 g 2.05x102g weight of element 100 100 Mole- molar mass of el

6. For Molybdenum, Weight-weight weight of ball 100 %x 0.238 x 3.756 g 100 = 8.9×10-3 g weight of element molar mass of eleme7. For Iron, weight %× weight of ball 9112 × 3756 g 3.67 g Weight = 100 100 weight of element Mole = molar mass of element 3.

3.7x10- mol 4.99 x 10- mol -0.74 3.74x10 mol Manganese = 4.99× 10-5 mol 9.37x10 mol 4.99×10-5 mol Carbon -18.7 = 19 9.27x10-

Let x be the number of atoms in each element removed. Since 12 sextillion atoms are removed, therefore lx + 6x + lx +7x+19x +